OpenStudy (anonymous):

a circle is tangent to the y axis at y=3 and has one x intercept at x=1. determine the other x intercept.

OpenStudy (amistre64):

has to be -1 i think

OpenStudy (amistre64):



OpenStudy (amistre64):

i think i seend it wrong

OpenStudy (amistre64):



OpenStudy (amistre64):

thats better

OpenStudy (amistre64):

x^2 + y^2 = r^2 0 + 3^2 = r^2 0 + 9 = r^2 r = 3

OpenStudy (amistre64):

and lets not forget (1,0) x^2 + y^2 = r^2 1 + 0 = r^2 taht theiry aint gonna fly

OpenStudy (amistre64):

(x-h)^2 + (y-k)^2 = r^2 (0-h)^2 + (3-k)^2 = r^2 (1-h)^2 + (0-k)^2 = r^2

OpenStudy (amistre64):

h^2 + 9 + k^2 -6k = r^2 1 + h^2 -2h + k^2 = r^2 h^2 + 9 + k^2 -6k = 1 + h^2 -2h + k^2 h^2-h^2 + 9-1 + k^2-k^2 -6k = -2h 8 -6k = -2h h = 3k -4

OpenStudy (amistre64):

maybe?

OpenStudy (amistre64):

i got it; the two point create a chord, whose midpoint at least is perped towards the center

OpenStudy (amistre64):

(0,3) -(1,0) ----- -1,-3 ; slope = 3, midpoint = (1/2, 3/2) and a slope of -1/3 when y = 3 and y = (-1/3)x +1/(2)3 + 3/2 are the same; we got the center

OpenStudy (amistre64):

3 = (-1/3)x +10/6 -3(3 - 10/6) = 3 -9 + 5 = x ..... missed something doh!! process works, math is off

OpenStudy (amistre64):

right at the beginning (0,3) -(1,0) ----- -1,3 ; slope = -3, midpoint = (1/2, 3/2) and a slope of 1/3 3 = (1/3)x -4/3 9 = x-4 13 = x the center should be at (13,3), which tells us that the other points is 13+12 away, or rather at (25,0) maybe

OpenStudy (amistre64):

soo close

OpenStudy (amistre64):

13,3 -: 1,0 -------- 12,3 = sqrt(144+9) = sqrt(123) sqrt(123) not= sqrt(169) were to short; lets try 12,3 to see where we go from here

OpenStudy (amistre64):

sqrt(121+9) = sqrt(130) not= sqrt(144) sqrt(100+9) = sqrt(109) not= sqrt(121) sqrt(81+9) = sqrt(90) not= sqrt(100) sqrt(64+9) = sqrt(73) not= sqrt(81) sqrt(49+9) = sqrt(58) not= sqrt(64) sqrt(36+9) = sqrt(45) not= sqrt(49) sqrt(25+9) = sqrt(34) not= sqrt(36) sqrt(16+9) = sqrt(25) = sqrt(25) thatll do it; center is at (5,3)

OpenStudy (amistre64):

1+5 = 5 + 4 = 9 9,0 and thats me final offer :) (x-5)^2 + (y-3)^2 = 25

OpenStudy (amistre64):

id have saved me self some trouble if i knew how to add :) \[3 = \frac{1}{3}x-\frac{1}{3(2)}+\frac{3}{2}\] \[9 = x-\frac{1}{(2)}+\frac{9}{2}\] \[18 = 2x-1+9\] \[18 = 2x+8\] \[10 = 2x\] \[5 = x\]

OpenStudy (anonymous):

The solution is incorrect. Could you please try again?

OpenStudy (amistre64):

which "solution" did you use?

OpenStudy (amistre64):

9,0 is the correct solution; once it is all said and done